OpenStudy (anonymous):

7+ log_2⁡〖y=3 log_2⁡y 〗

OpenStudy (anonymous):

\[7+\log_2(y)=3\log_2(2y)\]?

OpenStudy (anonymous):

Simplify 7+ log_2 y =3 log_2 y

OpenStudy (anonymous):

ok this is an equation, so probably the question is "solve" not "simplify"

OpenStudy (anonymous):

the y is on the same level as the log

OpenStudy (anonymous):

it says to simplify on the paper

OpenStudy (anonymous):

first subtract \[log_2(y)\] from both sides to get \[7=3\log_2(2y)-\log_2(y)\]

OpenStudy (anonymous):

no i'm sorry your right it is an equation, i was looking at the wrong question

OpenStudy (anonymous):

then use two properties of the log \[3\log_2(2y)=\log_2((2y)^3)=\log_2(8y^3)\]\]

OpenStudy (anonymous):

solve not simplify is right

OpenStudy (anonymous):

and also use \[\log_2(8y^3)-\log_2(y)=\log_2(\frac{8y^3}{y})\]

OpenStudy (anonymous):

no we are going to find y

OpenStudy (anonymous):

so now we have \[7=\log_2(8y^2)\]

OpenStudy (anonymous):

did you follow so far, because that is the important part

OpenStudy (anonymous):

no i'm lost, at the beginning when you subtracted from both sides

OpenStudy (anonymous):

oh ok lets go slow

OpenStudy (anonymous):

we want to solve for y

OpenStudy (anonymous):

so we are going to put everything with a y in it on one side, and the number on the other

OpenStudy (anonymous):

that is the reason for subtracting \[\log_2(y)\] from both sides ok?

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

so on the left side of the equal sign we have 7

OpenStudy (anonymous):

on the right side we have \[3log_2(2y)-\log_2(y)\]

OpenStudy (anonymous):

now our job is to write this as one log. that is write it as \[\log_2(\text{something})\]

OpenStudy (anonymous):

where did the (2y) come from

OpenStudy (anonymous):

oh crap i thought you wrote 2y but you did not

OpenStudy (anonymous):

sorry. the right side is just \[3\log_2(y)-\log_2(y)\] yes?

OpenStudy (anonymous):

oh ok, yes that's right

OpenStudy (anonymous):

whew sorry.

OpenStudy (anonymous):

it's fine

OpenStudy (anonymous):

now our job is still the same. to write it as a single log. that is write it as \[\log_2(\text{something})\]

OpenStudy (anonymous):

so we have \[3\log_2(y)=\log_2(y^3)\] is that step clear?

OpenStudy (anonymous):

we are using \[\log(x^n)=n\log(x)\] backwards

OpenStudy (anonymous):

clear or no?

OpenStudy (anonymous):

?how did you get the (y^3)

OpenStudy (anonymous):

ok it is a property of the logs that \[\log(x^n)=n\log(x)\]

OpenStudy (anonymous):

we had \[3\log_2(y)\]

OpenStudy (anonymous):

yes, yes. . .okay i get it, i remember learning that in class

OpenStudy (anonymous):

so i bring the 3 up as an exponent

OpenStudy (anonymous):

now we are going to use one more property of the log that says \[log(\frac{a}{b})=\log(a)-\log(b)\]

OpenStudy (anonymous):

have you seen that one?

OpenStudy (anonymous):

yes i have

OpenStudy (anonymous):

ok so the right hand side of the equal sign is now \[\log_2(y^3)-\log_2(y)\]

OpenStudy (anonymous):

which we rewrite as \[\log_2(\frac{y^3}{y})\]

OpenStudy (anonymous):

by that property. now we cancel to get \[\log_2(y^2)\]

OpenStudy (anonymous):

how we doing so far?

OpenStudy (anonymous):

because we are almost done!

OpenStudy (anonymous):

?what did you cancel

OpenStudy (anonymous):

i just canceled in \[\frac{y^3}{y}\] to get \[y^2\]

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

thats is i rewrote \[\log_2(\frac{y^3}{y})=\log_2(y^2)\]

OpenStudy (anonymous):

so finally (we are almost done, really) we have \[7=\log_2(y^2)\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

now to find y we rewrite this in "equivalent exponential form" that was the whole point of writing as a single log

OpenStudy (anonymous):

\[log_b(y)=x\] same as \[b^x=y\] so in this case we have \[\log_2(y^2)=7\] same as \[y^2=2^7\]

OpenStudy (anonymous):

making \[y=2^{\frac{7}{2}}\] or \[y=\sqrt{128}\] or even \[y=8\sqrt{2}\]

OpenStudy (anonymous):

btw not to confuse you , but when we had \[3\log_2(y)-\log_2(y)\] we could have written is as \[2\log_2(y)=\log_2(y^2)\] we get the same answer

OpenStudy (anonymous):

what does "btw" mean

OpenStudy (anonymous):

maybe simpler to subtract first and then use one property of the log

OpenStudy (anonymous):

by the way

OpenStudy (anonymous):

lol, oh okay

OpenStudy (anonymous):

what does lol mean?

OpenStudy (anonymous):

just kidding lol

OpenStudy (anonymous):

in any case we have the answer in many forms, whichever one you like

OpenStudy (anonymous):

lol, alright. . .thank you very much!

OpenStudy (anonymous):

yw